OpenStudy (mmend98):

How can 1/3x – 2 = y and 1/4x + 11 = y be set up as a system of equations? 3y – x = –6 4y – x = 44 3y + x = –6 4y + x = 44 3y – 3x = –6 4y – 4x = 44 3y + 3x = –6 4y + 4x = 44

OpenStudy (mmend98):

@ganeshie8

OpenStudy (mmend98):

@Michele_Laino

OpenStudy (anonymous):

convert them both to standard form.

OpenStudy (michele_laino):

for example, your first equation can be rewritten as follows: \[\frac{x}{3} - 2 = y \to x - 6 = 3y\] where I have multiplied both sides by 3

OpenStudy (mmend98):

and thats all? @Michele_Laino @dirtydishr4g

OpenStudy (michele_laino):

yes! Please multiply the both sides of your second equation by 4, what do you get?

OpenStudy (michele_laino):

Hint: you should get this: \[\frac{x}{4} + 11 = y \to x + 44 = 4y\]

OpenStudy (mmend98):

I get x+44=4y

OpenStudy (mmend98):

@Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

now, if I subtract x, I get: \[x + 44 - x = 4y - x\] please simplify

OpenStudy (mmend98):

4y-x=44? @Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

now I do the same with your first equation: \[x - 6 - x = 3y - x\] please simplify

OpenStudy (mmend98):

So the answer would be A? 3y – x = –6 4y – x = 44

OpenStudy (michele_laino):

that's right!

OpenStudy (mmend98):

Thank you. @Michele_Laino

OpenStudy (michele_laino):

:)